given endpoint (-6,2) and the midpoint (5,2
) find the other endpoint

Answer by stanbon(75887)

given endpoint (-6,2) and the midpoint (5,2)
find the other endpoint
-----
Fact: The mid-point coordinates are the average
of the corresponding end-point coordinates.
----------
Solve:
Let the end-point be (x,y)
(x+-6)/2 = 5
(y+2)/2 = 2
-----
x-6 = 10
y+2 = 4
---
x = 16
y = 2
-------------
Cheers,
Stan H.